The elastic stroke of the ping pong ball with a marble slab lasts delta t = 0.02 s. Calculate the impact force, knowing that the velocity of the strike ball was vo = 5 m / s.


1
Expert's answer
2021-01-31T03:39:30-0500

The ball weight for pin pong "2.7\ g=2.7\cdot 10^{-3}\ kg" .


"F=\frac{mv-mv_0}{\Delta t}=\frac{|2mv_0|}{\Delta t}=\frac{2\cdot2.7\cdot10^{-3}\cdot5}{0.02}=1.35\ (N)" . Answer